Hi Tutors:
Please show me how to solve this system of equations using the substitution method:
t=4-2s
t+2s=6
SAN ANTONIO truly appreciates help with Maths

Answer by longjonsilver(2297)

By substitution:

as t=4-2s, then we can say in eqn2:

(4-2s) + 2s = 6
4 - 2s + 2s = 6
which gives 4 = 6 which is patently not correct, so i am assuming you have written the question down wrong.

jon.